# FusionMamba: Dynamic Feature Enhancement for Multimodal Image Fusion with Mamba

### v1

@@ 4.1.2. Implementation Details and Metrics
Fusion performance is evaluated using six key metrics (Ronneberger et al., 2015; Vs et al., 2022): Structural Fidelity (SF), Structure Content Difference (SCD), Multiscale Structural Similarity Index Measure (MS-SSIM), Gradient-Based Metric ($\text{Q}^{AB/F}$ ), Feature Mutual Information (FMI), and Visual Information Fidelity (VIF)(Zhang and Demiris, 2023). These metrics evaluate fusion performance from multiple perspectives, including structural preservation, content difference, feature information, and visual fidelity in the fused image compared to source images. . SF measures structural fidelity, SCD quantifies content difference, MS-SSIM evaluates structural similarity across scales, $\text{Q}^{AB/F}$ assesses edge information retention, FMI measures feature information preservation, and VIF calculates visual fidelity comprehensively. These metrics together offer a comprehensive evaluation of fusion performance across different criteria.
@@ 4.1.3. Comparison Approaches
We perform extensive comparison experiments with state-of-the-art methods to demonstrate the validity of FusionMamba. These comparison methods include CSMCA (Xu et al., 2020b), U2Fusion (Guo et al., 2024), FusionGAN (Li et al., 2020), SDNet (Cvejic et al., 2005), MATR (Guan et al., 2023), IFT (Liu et al., 2024) and SwinFusion (Ma et al., 2020). CSMCA (Xu et al., 2020b) is the traditional fusion method. SDNet(Cvejic et al., 2005), FusionGAN, and U2Fusion (Guo et al., 2024) are CNN-based fusion methods. MATR(Guan et al., 2023), IFT and SwinFusion(Ma et al., 2020) are Transformer-based structures. We use the public codes with corresponding parameter settings.
@@
CSMCA (Xu et al., 2020b) 0.4745 1.1372 0.5986 30.2602 0.9444 0.8800
U2F (Guo et al., 2024) 0.4246 0.8399 0.2865 14.8600 0.8355 0.8619
FGAN (Li et al., 2020) 0.2597 0.1318 0.2068 12.1149 0.7234 0.8308
SDNet (Cvejic et al., 2005) 0.4224 0.9895 0.4797 28.9928 0.8672 0.8350
MATR (Guan et al., 2023) 0.4556 0.3468 0.5044 16.1920 0.6095 0.8600
SwinF (Ma et al., 2020) 0.6075 1.2659 0.6312 32.8764 0.9042 0.8339
IFT (Liu et al., 2024) 0.5016 1.1188 0.4571 21.4474 0.8928 0.8687
@@
CSMCA (Xu et al., 2020b) 0.5515 0.7892 0.6505 29.4565 0.9156 0.8401
U2F (Guo et al., 2024) 0.4618 1.2809 0.5153 26.3977 0.9074 0.8232
FGAN (Li et al., 2020) 0.3742 0.4620 0.2563 16.9172 0.6603 0.8127
SDNet (Cvejic et al., 2005) 0.5061 1.0270 0.5831 29.5060 0.8971 0.8208
MATR (Guan et al., 2023) 0.6033 0.6442 0.7254 26.4280 0.8290 0.8377
SwinF (Ma et al., 2020) 0.6775 1.3148 0.7254 32.2117 0.9261 0.8650
IFT (Liu et al., 2024) 0.6149 1.4068 0.5501 21.5242 0.9129 0.8431
@@ 4.2. Multimodal Medical Image Fusion
For the qualitative analysis, the first row in Fig. 6 displays the CT-MRI fusion task, including CT images, MRI images, and the fusion results of CSMCA (Xu et al., 2020b), U2Fusion (Guo et al., 2024), FusionGAN (Li et al., 2020), SDNet (Cvejic et al., 2005), MATR (Guan et al., 2023), SwinFusion (Ma et al., 2020), IFT (Liu et al., 2024), and the proposed FusionMamba. We zoomed in on the area in the red box, demonstrating that our fusion images can simultaneously retain the dense information of CT and the texture details of MRI. In Fig. 6, FusionGAN and U2Fusion (Guo et al., 2024) exhibit unsatisfactory brightness and clarity, with some gray matter blurring the texture details. Particularly, FusionGAN shows artifacts and redundant information. MATR (Guan et al., 2023)preserves structural details but weakens dense structures. SDNet(Cvejic et al., 2005) and CSMCA (Xu et al., 2020b) retain dense structures but lose some edge details. SwinFusion(Ma et al., 2020) achieves satisfactory fusion but with excessively sharp edges. In contrast, our FusionMamba retains more texture details while maintaining appropriate dense structures. Visually, they appear more natural with enhanced contrast.
The second row in Fig. 6 presents the PET-MRI fusion task. FusionGAN’s fusion images have lower contrast and missing texture details, while CSMCA (Xu et al., 2020b) and IFT(Liu et al., 2024) preserve texture information well. However, the darker color indicates insufficient preservation of functional information. Additionally, IFT(Liu et al., 2024), U2Fusion (Guo et al., 2024), and SDNet(Cvejic et al., 2005) retain functional features but with less sharp texture details in salient regions. SwinFusion(Ma et al., 2020) shows color distortion due to over-sharpening. In contrast, our FusionMamba preserves clear edges and texture details, with a color distribution more similar to PET images, enhancing visual perception.
@@
CSMCA (Xu et al., 2020b) 0.6051 0.1854 0.6301 13.708 0.9144 0.8827
U2F (Guo et al., 2024) 0.4525 0.5633 0.4789 12.7316 0.9209 0.8600
FGAN (Li et al., 2020) 0.3879 0.2734 0.3059 10.0998 0.8065 0.8484
SDNet (Cvejic et al., 2005) 0.4715 1.1937 0.4430 14.5479 0.8437 0.8785
MATR (Guan et al., 2023) 0.7235 0.1706 0.6620 12.6463 0.8883 0.8880
SwinF (Ma et al., 2020) 0.6147 1.1665 0.6468 15.7775 0.9568 0.8887
IFT (Liu et al., 2024) 0.6024 1.1504 0.6127 12.7114 0.9196 0.8770
@@ 4.3. Infrared and Visible Image Fusion
Figure 7 and Table 4 illustrate that RFN-Nest(Zhang and Ma, 2021), DATFuse, SDNet(Cvejic et al., 2005), and U2Fusion (Guo et al., 2024) struggle to accurately represent scene information in visible images due to inadequate global information interaction and improper intensity control. On the other hand, while IFT(Liu et al., 2024) and SwinFusion(Ma et al., 2020) manage to preserve some texture details in visible images, they are still susceptible to contamination from thermal radiation, resulting in a reduction in the clarity of infrared image targets to varying extents.
@@
RFN (Zhang and Ma, 2021) 0.612 1.6109 0.4229 11.3723 0.8544 0.8751
U2F (Guo et al., 2024) 0.4736 1.6390 0.3539 7.0651 0.8343 0.8836
FGAN (Li et al., 2020) 0.3476 1.2372 0.2107 5.6672 0.5314 0.6701
SDNet (Cvejic et al., 2005) 0.4715 1.6237 0.4430 14.5479 0.8436 0.8785
DATF (Ma et al., 2022) 0.6052 1.5297 0.5071 10.7585 0.7595 0.8730
SwinF (Ma et al., 2020) 0.6687 1.8095 0.5611 16.5580 0.9236 08894
IFT (Liu et al., 2024) 0.6350 1.7817 0.4668 11.0382 0.8606 0.8853
@@ 4.4. Multimodal Biomedical Image Fusion
To demonstrate the generalization capability of FusionMamba, we conduct green fluorescent protein (GFP) and phase contrast (PC) image fusion. GFP images provide functional information related to protein distribution, while PC images contain rich details of cell structure including nucleus and mitochondria. GFP and PC fusion images can promote biological research such as gene expression and protein function analysis. In Figure 8 and Table 5, it can be found that FusionGAN, U2Fusion (Guo et al., 2024) and IFT(Liu et al., 2024) can well preserve color information, but there is a certain loss of texture details. U2Fusion (Guo et al., 2024) and MATR(Guan et al., 2023) can well retain texture details, but there is slight color distortion. It is worth noting that both SwinFusion(Ma et al., 2020) and FusionMamba can effectively preserve texture and color information, which proves the effectiveness of cross fusion strategies in image fusion.
@@ 4.5. Computational Cost Analysis
The complexity evaluation in Table 6 assesses the operational efficiency of various methods by measuring floating-point operations per second (FLOPs) and running time. Specifically, the first image in the dataset undergoes testing within an infrared and visible light fusion scenario to compute the FLOPs of each method. We compare the CNN-based U2Fusion (Guo et al., 2024), IFCNN (Hwang et al., 2015) method with the Transformer-based IFT(Liu et al., 2024) and SwinFusion(Ma et al., 2020) method. The results reveal a notable advantage of the Mamba-based approach in terms of runtime, showcasing lower FLOPs and average uptime compared to CNN and Transformer methods. Importantly, our method demonstrates superior fusion performance when compared to mainstream image fusion algorithms.
@@
CSMCA (Xu et al., 2020b) 0.7609 1.4303 0.6475 11.1794 0.9453 0.8723
U2F (Guo et al., 2024) 0.6085 1.3442 0.3329 6.2236 0.8562 0.8616
FGAN (Li et al., 2020) 0.4538 1.2913 0.3027 8.6908 0.6823 0.8358
SDNet (Cvejic et al., 2005) 0.5046 1.3040 0.4983 14.0584 0.9023 0.8405
MATR (Guan et al., 2023) 0.8147 0.8400 0.7205 9.8661 0.8893 0.8637
SwinF (Ma et al., 2020) 0.8159 1.2185 0.6727 16.2299 0.9138 0.8612
IFT (Liu et al., 2024) 0.5652 1.3588 0.4659 12.0810 0.8993 0.8368
@@
Method U2F (Guo et al., 2024) IFCNN (Hwang et al., 2015) IFT (Liu et al., 2024) SwinF (Ma et al., 2020) FusionMamba (Ours)
@@ 5.1. Structure Ablation
To analyze the network architecture better, we conduct 5 sets of ablation experiments, as shown in Table 7. To validate DVSS effectiveness, we perform two additional experiments: EVSS and replacing DVSS with Transformer (Case 1 and Case 2). EVSS produces satisfactory results, but it’s less effective than DVSS in design due to DVSS’s stronger feature extraction capability. Replacing DVSS with the Transformer network leads to a decline in metrics. The proposed DFFM significantly enhances fusion effect, as seen in Case 3 where its absence causes a decrease in SF and $\mathrm{Q^{AB/F}}$. Removing the DFEM module (Case 4) shows similar results to not having DFFM. The Mamba module for cross-modal fusion is vital for effective information integration, as networks without Mamba module integration (Case 5) show degraded VIF task visual performance. Absence of CMFM leads to a decrease in MS-SSIM of fused images, indicating the importance of cross-domain integration for perceiving key goals and structures in fusion tasks.
@@ 6. Conclusion
In conclusion, our study addresses the challenges of multi-modal image fusion by proposing FusionMamba, a novel dynamic feature enhancement method integrated with the Mamba framework. Our approach combines an improved efficient Mamba model (EMamba) with dynamic convolution and channel attention, enhancing both global modeling capabilities and local feature extraction. We also introduce a dynamic feature fusion module (DFFM) that includes two dynamic feature enhancement modules (DFEM) and a cross-modality fusion Mamba module (CMFM), effectively enhancing texture, perception of differences, and correlation between modalities while suppressing redundant information. Our FusionMamba method has demonstrated state-of-the-art (SOTA) performance in various multimodal image fusion tasks. These results validate the generalization ability of our proposed method. For future work, we aim to investigate the application of FusionMamba in real-time scenarios and deploying it on resource-constrained devices would be beneficial for practical implementation. Furthermore, extending our evaluation to more diverse datasets and benchmarking against emerging fusion methods would provide a comprehensive understanding of the capabilities of FusionMamba.

### v2

@@ 4.1.2. Implementation Details and Metrics
Fusion performance is evaluated using six key metrics (Piella and Heijmans, 2003; Cvejic et al., 2005): Structural Fidelity (SF), Structure Content Difference (SCD), Multiscale Structural Similarity Index Measure (MS-SSIM), Gradient-Based Metric ($\text{Q}^{AB/F}$ ), Feature Mutual Information (FMI), and Visual Information Fidelity (VIF) (Chen and Varshney, 2007). These metrics evaluate fusion performance from multiple perspectives, including structural preservation, content difference, feature information, and visual fidelity in the fused image compared to source images. . SF measures structural fidelity, SCD quantifies content difference, MS-SSIM evaluates structural similarity across scales, $\text{Q}^{AB/F}$ assesses edge information retention, FMI measures feature information preservation, and VIF calculates visual fidelity comprehensively. These metrics together offer a comprehensive evaluation of fusion performance across different criteria.
@@ 4.1.3. Comparison Approaches
We perform extensive comparison experiments with state-of-the-art methods to demonstrate the validity of FusionMamba. These comparison methods include CSMCA (Liu et al., 2019), U2Fusion (Xu et al., 2020a), FusionGAN (Ma et al., 2019), SDNet (Zhang and Ma, 2021), MATR (Tang et al., 2022), IFT (Vs et al., 2022) and SwinFusion (Ma et al., 2022). CSMCA (Liu et al., 2019) is the traditional fusion method. SDNet (Zhang and Ma, 2021), FusionGAN, and U2Fusion (Xu et al., 2020a) are CNN-based fusion methods. MATR (Tang et al., 2022), IFT and SwinFusion (Ma et al., 2022) are Transformer-based structures. We use the public codes with corresponding parameter settings.
@@
CSMCA (Liu et al., 2019) 0.4745 1.1372 0.5986 30.2602 0.9444 0.8800
U2F (Xu et al., 2020a) 0.4246 0.8399 0.2865 14.8600 0.8355 0.8619
FGAN (Ma et al., 2019) 0.2597 0.1318 0.2068 12.1149 0.7234 0.8308
SDNet (Zhang and Ma, 2021) 0.4224 0.9895 0.4797 28.9928 0.8672 0.8350
MATR (Tang et al., 2022) 0.4556 0.3468 0.5044 16.1920 0.6095 0.8600
SwinF (Ma et al., 2022) 0.6075 1.2659 0.6312 32.8764 0.9042 0.8339
IFT (Vs et al., 2022) 0.5016 1.1188 0.4571 21.4474 0.8928 0.8687
@@
CSMCA (Liu et al., 2019) 0.5515 0.7892 0.6505 29.4565 0.9156 0.8401
U2F (Xu et al., 2020a) 0.4618 1.2809 0.5153 26.3977 0.9074 0.8232
FGAN (Ma et al., 2019) 0.3742 0.4620 0.2563 16.9172 0.6603 0.8127
SDNet (Zhang and Ma, 2021) 0.5061 1.0270 0.5831 29.5060 0.8971 0.8208
MATR (Tang et al., 2022) 0.6033 0.6442 0.7254 26.4280 0.8290 0.8377
SwinF (Ma et al., 2022) 0.6775 1.3148 0.7254 32.2117 0.9261 0.8650
IFT (Vs et al., 2022) 0.6149 1.4068 0.5501 21.5242 0.9129 0.8431
@@ 4.2. Multimodal Medical Image Fusion
For the qualitative analysis, the first row in Fig. 6 displays the CT-MRI fusion task, including CT images, MRI images, and the fusion results of CSMCA (Liu et al., 2019), U2Fusion (Xu et al., 2020a), FusionGAN (Ma et al., 2019), SDNet (Zhang and Ma, 2021), MATR (Tang et al., 2022), SwinFusion (Ma et al., 2022), IFT (Vs et al., 2022), and the proposed FusionMamba. We zoomed in on the area in the red box, demonstrating that our fusion images can simultaneously retain the dense information of CT and the texture details of MRI. In Fig. 6, FusionGAN and U2Fusion (Xu et al., 2020a) exhibit unsatisfactory brightness and clarity, with some gray matter blurring the texture details. Particularly, FusionGAN shows artifacts and redundant information. MATR (Tang et al., 2022)preserves structural details but weakens dense structures. SDNet (Zhang and Ma, 2021) and CSMCA (Liu et al., 2019) retain dense structures but lose some edge details. SwinFusion (Ma et al., 2022) achieves satisfactory fusion but with excessively sharp edges. In contrast, our FusionMamba retains more texture details while maintaining appropriate dense structures. Visually, they appear more natural with enhanced contrast.
The second row in Fig. 6 presents the PET-MRI fusion task. FusionGAN’s fusion images have lower contrast and missing texture details, while CSMCA (Liu et al., 2019) and IFT (Vs et al., 2022) preserve texture information well. However, the darker color indicates insufficient preservation of functional information. Additionally, IFT (Vs et al., 2022), U2Fusion (Xu et al., 2020a), and SDNet (Zhang and Ma, 2021) retain functional features but with less sharp texture details in salient regions. SwinFusion (Ma et al., 2022) shows color distortion due to over-sharpening. In contrast, our FusionMamba preserves clear edges and texture details, with a color distribution more similar to PET images, enhancing visual perception.
@@
CSMCA (Liu et al., 2019) 0.6051 0.1854 0.6301 13.708 0.9144 0.8827
U2F (Xu et al., 2020a) 0.4525 0.5633 0.4789 12.7316 0.9209 0.8600
FGAN (Ma et al., 2019) 0.3879 0.2734 0.3059 10.0998 0.8065 0.8484
SDNet (Zhang and Ma, 2021) 0.4715 1.1937 0.4430 14.5479 0.8437 0.8785
MATR (Tang et al., 2022) 0.7235 0.1706 0.6620 12.6463 0.8883 0.8880
SwinF (Ma et al., 2022) 0.6147 1.1665 0.6468 15.7775 0.9568 0.8887
IFT (Vs et al., 2022) 0.6024 1.1504 0.6127 12.7114 0.9196 0.8770
@@ 4.3. Infrared and Visible Image Fusion
Fig. 7 and Table 4 illustrate that RFN-Nest(Li et al., 2021), DATFuse (Tang et al., 2023a), SDNet (Zhang and Ma, 2021), and U2Fusion (Xu et al., 2020a) struggle to accurately represent scene information in visible images due to inadequate global information interaction and improper intensity control. On the other hand, while IFT (Vs et al., 2022) and SwinFusion (Ma et al., 2022) manage to preserve some texture details in visible images, they are still susceptible to contamination from thermal radiation, resulting in a reduction in the clarity of infrared image targets to varying extents.
@@
RFN (Li et al., 2021) 0.612 1.6109 0.4229 11.3723 0.8544 0.8751
U2F (Xu et al., 2020a) 0.4736 1.6390 0.3539 7.0651 0.8343 0.8836
FGAN (Ma et al., 2019) 0.3476 1.2372 0.2107 5.6672 0.5314 0.6701
SDNet (Zhang and Ma, 2021) 0.4715 1.6237 0.4430 14.5479 0.8436 0.8785
DATF (Tang et al., 2023a) 0.6052 1.5297 0.5071 10.7585 0.7595 0.8730
SwinF (Ma et al., 2022) 0.6687 1.8095 0.5611 16.5580 0.9236 08894
IFT (Vs et al., 2022) 0.6350 1.7817 0.4668 11.0382 0.8606 0.8853
@@ 4.4. Multimodal Biomedical Image Fusion
To demonstrate the generalization capability of FusionMamba, we conduct green fluorescent protein (GFP) and phase contrast (PC) image fusion. GFP images provide functional information related to protein distribution, while PC images contain rich details of cell structure including nucleus and mitochondria. GFP and PC fusion images can promote biological research such as gene expression and protein function analysis. In Figure 8 and Table 5, it can be found that FusionGAN (Ma et al., 2019), U2Fusion (Xu et al., 2020a) and IFT (Vs et al., 2022) can well preserve color information, but there is a certain loss of texture details. U2Fusion (Xu et al., 2020a) and MATR (Tang et al., 2022) can well retain texture details, but there is slight color distortion. It is worth noting that both SwinFusion (Ma et al., 2022) and FusionMamba can effectively preserve texture and color information, which proves the effectiveness of cross fusion strategies in image fusion.
@@ 4.5. Computational Cost Analysis
The complexity evaluation in Table 6 assesses the operational efficiency of various methods by measuring floating-point operations per second (FLOPs) and running time. Specifically, the first image in the dataset undergoes testing within an infrared and visible light fusion scenario to compute the FLOPs of each method. We compare the CNN-based U2Fusion (Xu et al., 2020a), IFCNN (Zhang et al., 2020) method with the Transformer-based IFT (Vs et al., 2022) and SwinFusion (Ma et al., 2022) method. The results reveal a notable advantage of the Mamba-based approach in terms of runtime, showcasing lower FLOPs and average uptime compared to CNN and Transformer methods. Importantly, our method demonstrates superior fusion performance when compared to mainstream image fusion algorithms.
@@
CSMCA (Liu et al., 2019) 0.7609 1.4303 0.6475 11.1794 0.9453 0.8723
U2F (Xu et al., 2020a) 0.6085 1.3442 0.3329 6.2236 0.8562 0.8616
FGAN (Ma et al., 2019) 0.4538 1.2913 0.3027 8.6908 0.6823 0.8358
SDNet (Zhang and Ma, 2021) 0.5046 1.3040 0.4983 14.0584 0.9023 0.8405
MATR (Tang et al., 2022) 0.8147 0.8400 0.7205 9.8661 0.8893 0.8637
SwinF (Ma et al., 2022) 0.8159 1.2185 0.6727 16.2299 0.9138 0.8612
IFT (Vs et al., 2022) 0.5652 1.3588 0.4659 12.0810 0.8993 0.8368
@@
Method U2F (Xu et al., 2020a) IFCNN (Zhang et al., 2020) IFT (Vs et al., 2022) SwinF (Ma et al., 2022) FusionMamba (Ours)
@@ 5.1. Structure Ablation
To analyze the network architecture better, we conduct 5 sets of ablation experiments, as shown in Table 7. To validate DVSS effectiveness, we perform two additional experiments: EVSS (Pei et al., 2024) and replacing DVSS with Transformer (Case 1 and Case 2). EVSS produces satisfactory results, but it’s less effective than DVSS in design due to DVSS’s stronger feature extraction capability. Replacing DVSS with the Transformer network leads to a decline in metrics. The proposed DFFM significantly enhances fusion effect, as seen in Case 3 where its absence causes a decrease in SF and $\mathrm{Q^{AB/F}}$. Removing the DFEM module (Case 4) shows similar results to not having DFFM. The Mamba module for cross-modal fusion is vital for effective information integration, as networks without Mamba module integration (Case 5) show degraded VIF task visual performance. Absence of CMFM leads to a decrease in MS-SSIM of fused images, indicating the importance of cross-domain integration for perceiving key goals and structures in fusion tasks.
@@ 6. Conclusion
In conclusion, our study addresses the challenges of multi-modal image fusion by proposing FusionMamba, a novel dynamic feature enhancement method integrated with the Mamba framework. Our approach combines an improved efficient Mamba model with dynamic convolution and channel attention, enhancing both global modeling capabilities and local feature extraction. We also introduce a dynamic feature fusion module (DFFM) that includes two dynamic feature enhancement modules (DFEM) and a cross-modality fusion Mamba module (CMFM), effectively enhancing texture, perception of differences, and correlation between modalities while suppressing redundant information. Our FusionMamba method has demonstrated state-of-the-art (SOTA) performance in various multimodal image fusion tasks. These results validate the generalization ability of our proposed method. For future work, we aim to investigate the application of FusionMamba in real-time scenarios and deploying it on resource-constrained devices would be beneficial for practical implementation. Furthermore, extending our evaluation to more diverse datasets and benchmarking against emerging fusion methods would provide a comprehensive understanding of the capabilities of FusionMamba.